A box contains nickels, dimes and pennies worth $6.20. The number of nickels is two less than twice the number of dimes and there is an equal number of pennies and dimes. How many nickels are in the box?
Answer by lwsshak3(11628)

A box contains nickels, dimes and pennies worth $6.20. The number of nickels is two less than twice the number of dimes and there is an equal number of pennies and dimes. How many nickels are in the box?
***
let x=number of pennies and dimes
2x-2=number of nickels
..
.10x+.01x+.05(2x-2)=6.20
.10x+.01x+.10x-.10=6.20
.21x6.30
x=30
2x-2=58
How many nickels are in the box? 58